# On the use of Graphs for Satellite Image Time Series

### v1

@@ Abstract
Keywords: Spatio-temporal graph, satellite image time series, graph theory, graph neural network, multitemporal imagery, object-based image analysis, remote sensing. This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
The Earth is constantly evolving. Its surface follows complex dynamics orchestrated by a wide variety of phenomena acting on all scales, from global movements of tectonic plates, atmosphere and oceans to local changes linked to the life cycles of forests, agriculture or human impacts. Observing the Earth from space enables global studies that are not possible with in-situ measurements, both in terms of spatial coverage and temporal sampling frequency. According to the Satellite Database maintained by the Union of Concerned Scientists [1], there were more than one thousand operational satellites dedicated to Earth observation in May 2023. These instruments capture satellite images of any location in the world at high spatial and temporal resolutions, enabling to study natural and artificial phenomena on the Earth’s surface.
@@ 1.1 From object-based to graph-based image analysis
Indeed, SITS can be voluminous spatially, temporally and spectrally, depending on their resolution in each of these dimensions, leading to a high variability of recorded measurements and a memory-costly storage and processing. By processing datacubes at a pixel scale, most of the existing techniques [3] struggle to handle these problems. Historically, similar issues have arisen with high-resolution satellite images. Beyond computational complexity, partially solved by hardware [4] and data management and distribution [5, 6, 7, 8] progress, the Object-Based Image Analysis (OBIA) paradigm was also used [9, 10] to overcome these limitations. This framework extends the consideration of features to spatial areas rather than independent pixels, by abandoning the regular Euclidean structure of satellite data (i.e., the pixel grid) and considering objects as working units. Moreover, these objects attempt to provide a faithful real-world representation while limiting the data volume, and so assumptions are made about their inherent distribution, for example by defining a homogeneity criterion for reflectances within regions. In the OBIA paradigm, an object is the representation in an information system of a real-world phenomenon, that may exist as a physical entity, like a building or river, or a more abstract concept, such as a state or country. Therefore, OBIA enables adding semantics to the data, facilitating image understanding by distinguishing meaningful entities and scene components. In short, an OBIA approach identifies objects, usually through segmentation, extracts object features, and finally analyzes these objects on the basis of rule-based algorithms or machine learning.
@@
OBIA tries to mimic some aspects of human vision such as spatial concepts that group local elements, i.e., pixels, to identify real objects as geographical regions. To describe these objects, most of the OBIA methods use inherent object attributes such as spectral, textural and geometrical features. However, in some applications, an object is better described by its context and the other surrounding objects (e.g., islands, urban parks), or by its inner components and the relations between them (e.g., residential areas). Therefore incorporating a structural approach with object relationships to OBIA seems primordial [11]. Existing OBIA works that take into account an object’s environment does so by extracting new features through the use of multi-scale extraction [12, 13] or object hierarchies [14], for example. However, such contextual features do not exploit the inner characteristics of other objects, preferring to analyze object’s context from scratch, which limits the representation of relationships, especially the long-range ones. To overcome this issue, a natural and flexible way of modeling relationships between objects is to structure data with a graph.
@@
The graph-based approach to model satellite image content as spatial objects interacting with each other can be considered as an extension of the OBIA paradigm that explicitly adds and exploits inter-object relationships. It retains the advantages of OBIA for representing scenes as sets of geographic objects, but also allows the mathematical tools of graph theory to be used to analyze them. In this context, the definition of an object is extended beyond the purely semantic sense to include any pixel sets aggregated together by an arbitrary criterion. In the most granular case, even individual pixels can be treated as objects. In the remainder of the paper, we will consider this generic definition of an object. Graph representation is memory efficient, suitable not only for computations, but also for human interpretation, thanks to its powerful synthesis and visualization capabilities [15]. Graphs are highly flexible mathematical tools that can be used to model all kinds of local contextual interactions, e.g., geographical relationships, but also to take into account long-term connections, semantic relationships and experts’ prior knowledge of the domain [16].
1.2 Graph-based satellite image time series
@@
(c) Spatio-temporal graph
(d) Visualization
(e) Pattern mining
(f) Classification
(g) Forecasting
@@
Figure 1: Pipeline of a spatio-temporal graph used as an intermediate representation of satellite image time series (SITS) for different tasks. (a) SITS enable the study of phenomena occurring at the Earth’s surface (Section 3.1). (b) The entities involved in the captured phenomena are identified and modeled in a spatio-temporal graph with their spatial and spatio-temporal relationships (Section 3.2). The plot shows entities as black nodes, spatial edges as grey lines and spatio-temporal edges as blue dotted lines. This representation can then be exploited for the targeted application. (c) Graphs can serve as a visualization tool for experts (Section 4.1). The plot illustrates the normalized difference water index (NDWI) temporal profile of the entities, in the same way as in [29], and suggests a soil covering event by water. (d) Graphs can help to detect predefined or frequent patterns (Section 4.2). In the illustration, merger and continuation patterns, as defined in [30], are highlighted. (e) Graphs can be used for classification tasks in time and space (Section 4.3). (f) Graphs can efficiently handle the way entities dynamically interact with each other and thus provide anticipation of the future state of a component of the Earth system as part of a forecasting task (Section 4.4).
@@
Beyond SITS modeling, graphs have a rich history of applications, including computational chemistry, recommendation systems, image processing, traffic forecasting and physics simulation [31]. Among these applications, graphs can serve as a powerful tool for interpreting large volumes of data. The work of Sun et al. [16] provides a survey of remote sensing image interpretation with semantic graph-based methods. It explores the different ways to represent remote sensing knowledge and how graph-based techniques are used for remote sensing image understanding. Although with a different focus than ours, this study shows the added value that graphs bring to the extraction of interpretable knowledge in the remote sensing field. This graph data processing requires specific techniques that have evolved considerably over the years [32, 33]. In fact, nowadays, increasing attention is being paid to graph neural networks (GNNs) and their transfer to the field of remote sensing [34, 35]. A recent review offers an overview of GNNs applied to remote sensing and proposes two taxonomies based on either input data or tasks [36]. It identifies the benefits of deep graph learning in remote sensing, while pointing to future directions for improving the interaction between these two fields. The taxonomy based on input data proposed in [36] mentions spatio-temporal data, but only a couple of works using SITS are actually cited [37, 38] even though the use of graphs to represent such data predates the emergence of deep learning. In fact, graphs can be used to handle and detect changes from bi-temporal satellite imagery [39, 40], which leads to considering an extension to the multi-temporal case of SITS [2]. The development of spatio-temporal graphs for Earth observation data, including SITS, is also recommended for future research by Zhao et al. [35], which provide the readers with a qualitative review on GNNs for Earth observation, covering a broad spectrum of application fields and data sources.
@@
These studies have highlighted the benefits of graph structure as a powerful tool for representing and analyzing remote sensing data. A focus on spatio-temporal data is lacking, however, despite the interest in such graph-based approaches. In fact, creating a graph from SITS is also complex, due to the wide freedom of assumptions and possible interactions between the objects. Thus, our work aims to examine the use of graphs to exploit SITS, from the creation of the graph and design choices, to graph analysis applied to downstream tasks. It is motivated by the need to bridge the gap between traditional graph-based methods applied to remote sensing and the recent interest given to this mathematical tool since the deep learning era. In addition, the observed entities are constantly interacting with each other, which makes it promising to exploit spatio-temporal relationships with graphs. Our study provides a broader view of graphs specifically used for satellite-based spatio-temporal data, investigating both graph extraction from SITS and traditional analysis methods, but also current trends in deep learning with GNNs. Figure 1 illustrates the whole graph-based pipeline studied in this paper to process SITS. With a comprehensive review and two case studies, this paper will hopefully provide the remote sensing and graph learning communities with guidance to overcome the challenges of creating and operating SITS using graphs. To stimulate cross-domain interactions, we guide this paper through three research questions:
@@ 2 Background on Graph Theory
Figure 2: Examples to illustrate graph terminology: (a) An undirected graph with node and edge attributes. The green shaded area, denoted by $\mathcal{N}(\mathrm{v}_{5})$, depicts the neighborhood of node $\mathrm{v}_{5}$. (b) A multi-relational graph with weighted directed edges. The edge width represents the edge weight and orange and blue colored-edges represents two types of relationship.
@@ 2.1 Graph definition
It is worth noting that additional information can be associated with the graph. For example, an attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X}_{\mathcal{V}},\mathcal{X}_{\mathcal{E}},\mathcal{X}_{\mathcal{G}})$ is one where each node, each edge and the graph itself can be associated with specific properties that describe its distinctive features, respectively represented by $\mathcal{X}_{\mathcal{V}}=\left\{\mathrm{x}_{i}\mid\mathrm{v}_{i}\in\mathcal{V}\right\}\in\mathbb{R}^{|\mathcal{V}|\times f_{\mathcal{V}}}$, $\mathcal{X}_{\mathcal{E}}=\left\{\mathrm{x}_{ij}\mid\mathrm{e}_{ij}\in\mathcal{E}\right\}\in\mathbb{R}^{|\mathcal{E}|\times f_{\mathcal{E}}}$ and $\mathcal{X}_{\mathcal{G}}\in\mathbb{R}^{f_{\mathcal{G}}}$, with $f_{\mathcal{V}}$, $f_{\mathcal{E}}$ and $f_{\mathcal{G}}$ the dimensions of feature vectors. A graph can also have weights on its edges to model the strength of interaction and is then referred as a weighted graph, with $w_{ij}\in\mathbb{R}$ being the weight of $e_{ij}$. The edges can also represent different kinds of relationships with a multi-relational graph having $\mathcal{E}=\mathcal{E}^{1}\sqcup\mathcal{E}^{2}\sqcup\cdots\sqcup\mathcal{E}^{r}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$, where $\mathcal{R}$ is a set of $r$ relationship types and $\sqcup$ is the disjoint union operator. To go further than its topological connections and features, a graph may also be endowed with some spatial geometry, being thus a geometric graph [43]. The geometrical information $\mathcal{X}_{P_{\mathcal{V}}}=\left\{\mathrm{p}_{i}\mid\mathrm{v}_{i}\in\mathcal{V}\right\}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is usually carried by nodes, in addition to other features, on a $d$-dimensional manifold. The most common geometrical feature example for remote sensing data is the spatial coordinates. These extended definitions can obviously be combined together according to the degree of realism and complexity required. Figure 2 shows examples of graphs. In particular, it represents an attributed undirected graph and a multi-relational graph with weighted directed edges.
@@ 3.1 Collecting the satellite image time series
In this section, the focus is given to the spatial and temporal characteristics, overlooking data storage and transmission constraints, and preprocessing such as radiometric correction [58] and geometric realignment [59]. Specifically, spatial resolution and revisit time, i.e., the time it takes for the satellite to acquire new data from the same location on Earth, depend mainly on the satellite’s orbit and swath.
Geostationary satellites enable high temporal sampling, generally less than an hour, of a large, defined area. However, due to the distance of the orbit and the limited capacity of imaging instruments, spatial resolution is at best hectometric. Weather satellites like Meteosat Third Generation (MTG) and Geostationary Operational Environmental Satellites (GOES) generally use this orbit [60, 61].
For applications requiring higher spatial resolution and global coverage, low earth orbits, often sun-synchronous, are used. For example, these orbits are used by the Landsat program which is the first civilian Earth observation space program and provides the longest continuous space-based record of Earth’s land starting from 1972 [62]. However, these satellites have a limited field of view, enabling them to observe only a part of the Earth at any given time, so their revisit time at nadir is at best a couple of weeks. To increase temporal resolution, a wider swath can be used, allowing significant overlap between orbits and more frequent acquisitions of the same location. But in general, a wider swath tends to reduce spatial resolution. MODIS and PROBA-V sensors use this principle to study landscape and vegetation every one to two days but cannot capture the finest details due to spatial resolution ranging from a few hundred meters to a kilometer [63, 64].
Another technique for reducing the revisit time of the same area is using a second identical satellite on the same orbit as the first, but delayed. ESA’s Sentinel missions are employing this approach increasing the resolution to a few meters for weekly acquisitions [65]. Adding more satellites to the constellations obviously improves temporal resolution without compromising spatial resolution. New generations of satellites are being deployed in mega-constellations, such as the SkySat and PlanetScope constellations [66]. The PlanetScope constellation includes over a hundred satellites, allowing a daily revisit time with a three meters ground sample distance, but raising the question of low-Earth orbit congestion [67]. The SkySat satellites enables sub-metric acquisitions thanks to its narrow swath width with a revisit capability of specific sites of five days, which is increased up to seven times a day thanks to the 21 satellites of the constellation.
In fact, the use of agile satellites that are able to move away from nadir for acquisition with steerable mirrors, such as SPOT 1-5, or gyroscopic controllers, such as Pleiades and Pleiades Neo, reduces revisit time. They are capable of maintaining very high spatial resolution while still being able to image specifics land surface plots frequently. The constraint lies mainly in the scheduling of acquisitions, the data storage and transfer capabilities, as repeated acquisition of a given site must be planned and may compete with other requests [68]. This capability of revisiting targets of interest is however different to the revisit time for a global coverage, and is specific to agile satellites. They are thus suited for revisiting regions of interest from time to time. The SITS obtained from these images are much more complex to process, due to the multiple viewing angles and different positions of the sun, which result in cast shadows. However, this does not prevent the identification of objects and the use of graphs to model their spatial and temporal relationships.
@@
These differences in acquisition methods mean that SITS have a wide range of spatial resolutions, from a few decimeters to several kilometers, and different temporal samplings, regular or not. There is a third concern regarding data resolution. For optical SITS, it corresponds to spectral resolution ranging from panchromatic acquisition, to multispectral with a few wavelengths, often in the visible and infrared range, and hyperspectral with acquisitions over hundred of narrowed contiguous spectral bands. It should be noted that there is a trade-off between spatial and spectral resolutions, meaning that a system with high spectral resolution can only provide medium or low spatial resolution. For radar data, we can consider the number of frequency bands and the polarization of the acquisition as the radar counterpart of spectral resolution. The ability to distinguish subtle differences in measured energies is called radiometric resolution. For more information on remote sensing platforms and sensors, see [57, 69]. A lot of information about specific satellites is also available on eoPortal (www.eoportal.org).
@@
The analysis of SITS is likely to extend beyond the multispectral and radar aspects currently studied. For example, some works manipulate multitemporal hyperspectral images—such as those provided by EnMAP, EO-1 Hyperio, or the future CHIME mission—in the form of time series at the pixel level. However, only a few exploit these satellite acquisitions as SITS, exploiting jointly spatial, temporal and spectral dimensions [70, 71, 72, 73, 74], certainly due to the lack of methods to process such complex and large data. Moreover, if we consider future satellite planning, the CO3D constellation will add depth to the spatial dimension thanks to three-dimensional acquisitions, going beyond the stereo acquisitions of Pléiades and Pléiades Neo, and thus providing multi-temporal 3D point clouds [75], which can be processed using graph-based methods.

### v2

# arXiv:2505.16685v2[cs.CV] 27 Feb 2026

@@ Abstract
Keywords: Spatio-temporal graph, satellite image time series, graph theory, graph neural network, multitemporal imagery, object-based image analysis, remote sensing. DOI: 10.1109/MGRS.2025.3622200
2473-2397 © 2025 IEEE. All rights reserved, including rights for text and data training of artificial intelligence and similar technologies.
@@ 1 Introduction
The Earth is constantly evolving. Its surface follows complex dynamics orchestrated by a wide variety of phenomena acting on all scales, from global movements of tectonic plates, atmosphere and oceans to local changes linked to the life cycles of forests, agriculture or human impacts. Observing the Earth from space enables global studies that are not possible with in-situ measurements, both in terms of spatial coverage and temporal sampling frequency, as satellites capture images of any location in the world at high spatial and temporal resolutions, enabling the study of natural and artificial phenomena on the Earth’s surface. According to the Satellite Database maintained by the Union of Concerned Scientists [64], there were more than one thousand operational satellites dedicated to Earth observation in May 2023.
@@ 1.1 From object-based to graph-based image analysis
Indeed, SITS can be voluminous spatially, temporally and spectrally, depending on their resolution in each of these dimensions, leading to a high variability of recorded measurements and a memory-costly storage and processing. By processing datacubes at a pixel scale, most of the existing techniques [14] struggle to handle these problems. Historically, similar issues have arisen with high-resolution satellite images. Beyond computational complexity, partially solved by hardware [34] and data management and distribution [132, 189, 63, 104] progress, the Object-Based Image Analysis (OBIA) paradigm was used [13, 111] to overcome these limitations. This framework extends the consideration of features to spatial areas rather than independent pixels, by abandoning the regular Euclidean structure of satellite data (i.e., the pixel grid) and considering objects as working units.
@@
OBIA tries to mimic some aspects of human vision such as spatial concepts that group local elements, i.e., pixels, to identify real objects as geographical regions, like a river, a building or a field. To describe these objects, most of the OBIA methods use inherent object attributes such as spectral, textural and geometrical features. However, in some applications, an object is better described by its context and the other surrounding objects (e.g., islands, urban parks), or by its inner components and the relations between them (e.g., residential areas). Therefore incorporating a structural approach with object relationships to OBIA seems primordial [194]. To overcome this issue, a natural and flexible way of modeling relationships between objects is to structure data with a graph.
@@
The graph-based approach retains the advantages of OBIA for representing scenes as sets of geographic objects, but also allows the mathematical tools of graph theory to be used to analyze them. In this context, the definition of an object is extended beyond the purely semantic sense to include any pixel sets aggregated together by an arbitrary criterion. In the most granular case, even individual pixels can be treated as objects. In the remainder of the paper, we will consider this generic definition of an object. Graph representation is memory efficient, suitable not only for computations, but also for human interpretation, thanks to its powerful synthesis and visualization capabilities [204]. Graphs are highly flexible mathematical tools that can be used to model all kinds of local contextual interactions, e.g., geographical relationships, but also to take into account long-term connections, semantic relationships and experts’ prior knowledge of the domain [190].
@@
(c) Spatio-temporal objects
(e) Spatio-temporal graph
(f) Visualization
(g) Pattern mining
(h) Classification
(i) Forecasting
@@
Figure 1: Pipeline of a spatio-temporal graph used as an intermediate representation of satellite image time series (SITS) for different tasks. (a) SITS enable the study of phenomena occurring at the Earth’s surface (Section 3.1). (b) The entities involved in the captured phenomena are identified, modeled as a set of spatio-temporal objects and outlined in red (Section 3.2.1). (c) The model is completed by spatial and spatio-temporal relationships between entities, forming a spatio-temporal graph (Section 3.2.2). The plot shows entities as black nodes, spatial edges as grey lines and spatio-temporal edges as blue dotted lines. This representation can then be exploited for the targeted application. (d) Graphs can serve as a visualization tool for experts (Section 4.1). The plot illustrates the normalized difference water index (NDWI) temporal profile of the entities, in the same way as in [70], and suggests a soil covering event by water. (e) Graphs can help to detect predefined or frequent patterns (Section 4.2). In the illustration, merger and continuation patterns, as defined in [212], are highlighted. (f) Graphs can be used for classification tasks in time and space (Section 4.3). (g) Graphs can efficiently handle the way entities dynamically interact with each other and thus provide anticipation of the future state of a component of the Earth system as part of a forecasting task (Section 4.4).
1.2 Related surveys and Research questions
@@
Beyond SITS modeling, graphs have a rich history of applications, including computational chemistry, recommendation systems, image processing, traffic forecasting and physics simulation [151]. Among these applications, graphs can serve as a powerful tool for interpreting large volumes of data, especially with specific graph data processing techniques that have evolved considerably over the years [75, 24]. In the field of image processing, graphs have been established as a natural representation for modeling spatial relationships and structures, as highlighted in the seminal book Image Processing and Analysis with Graphs [116]. Extending this perspective to Earth observation, graph-based techniques have also been employed to represent and analyze remote sensing knowledge, as reviewed by Sun et al. [190]. In fact, nowadays, increasing attention is being paid to graph neural networks (GNNs) and their transfer to the field of remote sensing [130, 228]. A recent review offers an overview of GNNs applied to remote sensing and proposes two taxonomies based on either input data or tasks [103]. It identifies the benefits of deep graph learning in remote sensing, while pointing to future directions for improving the interaction between these two fields. The development of spatio-temporal graphs for Earth observation data, including SITS, is recommended for future research by Zhao et al. [228], who provide the readers with a qualitative review on GNNs for Earth observation, covering a broad spectrum of application fields and data sources. In contrast, another recent review on SITS [139] highlights the potential of graphs for SITS analysis, but provides only a preliminary overview.
@@
These studies have highlighted the benefits of graph structure as a powerful tool for representing and analyzing images, and more specifically remote sensing data. A focus on spatio-temporal data is lacking, despite the interest in such graph-based approaches. In fact, creating a graph from SITS is also complex, due to the wide freedom of assumptions and possible interactions between the objects. Thus, our study provides a broader view of graphs specifically used for satellite-based spatio-temporal data, investigating both graph extraction from SITS and analysis methods, including current trends in deep graph learning. Figure 1 illustrates the whole graph-based pipeline studied in this paper to process SITS. With a comprehensive review and two case studies, this paper will hopefully provide the remote sensing and graph learning communities with guidance to overcome the challenges of creating and operating SITS using graphs. To stimulate cross-domain interactions, we guide this paper through three research questions:
@@ 2 Background on Graph Theory
Figure 2: Examples to illustrate graph terminology. (a) An undirected graph with node and edge attributes. The green shaded area, denoted by $\mathcal{N}(\mathrm{v}_{5})$, depicts the neighborhood of node $\mathrm{v}_{5}$. (b) A multi-relational graph with weighted directed edges. The edge width represents the edge weight and orange and blue colored-edges represents two types of relationship.
@@ 2.1 Graph definition
It is worth noting that additional information can be associated with the graph. For example, an attributed graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{X}_{\mathcal{V}},\mathcal{X}_{\mathcal{E}},\mathcal{X}_{\mathcal{G}})$ is one where each node, each edge and the graph itself can be associated with specific properties that describe its distinctive features, respectively represented by $\mathcal{X}_{\mathcal{V}}=\left\{\mathbf{x}_{i}\mid\mathrm{v}_{i}\in\mathcal{V}\right\}\in\mathbb{R}^{|\mathcal{V}|\times f_{\mathcal{V}}}$, $\mathcal{X}_{\mathcal{E}}=\left\{\mathbf{x}_{ij}\mid\mathrm{e}_{ij}\in\mathcal{E}\right\}\in\mathbb{R}^{|\mathcal{E}|\times f_{\mathcal{E}}}$ and $\mathcal{X}_{\mathcal{G}}\in\mathbb{R}^{f_{\mathcal{G}}}$, with $_{\mathcal{V}}, ff_{\mathcal{E}}$ and $f_{\mathcal{G}}$ the dimensions of feature vectors. A graph can also have weights on its edges to model the strength of interaction and is then referred as a weighted graph, with $w_{ij}\in\mathbb{R}$ being the weight of $\mathrm{e}_{ij}$. The edges can also represent different kinds of relationships with a multi-relational graph having $\mathcal{E}=\mathcal{E}^{1}\sqcup\mathcal{E}^{2}\sqcup\cdots\sqcup\mathcal{E}^{r}\subseteq\mathcal{V}\times\mathcal{R}\times\mathcal{V}$, where $\mathcal{R}$ is a set of $r$ relationship types and $\sqcup$ is the disjoint union operator. To go further than its topological connections and features, a graph may also be endowed with some spatial geometry, being thus a geometric graph [21]. The geometrical information $\mathcal{X}_{P_{\mathcal{V}}}=\left\{\mathbf{p}_{i}\mid\mathrm{v}_{i}\in\mathcal{V}\right\}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is usually carried by nodes, in addition to other features, on a $d$-dimensional manifold. The most common geometrical feature example for remote sensing data is the spatial coordinates. These extended definitions can obviously be combined together according to the degree of realism and complexity required. Figure 2 shows examples of graphs. In particular, it represents an attributed undirected graph and a multi-relational graph with weighted directed edges.
@@ 2.2 Spatio-temporal graph
In the graph literature, there are other mathematical formulations for describing a dynamic phenomenon that changes over time. The most used are the Discrete-Time Dynamic Graph (DTDG) and the Continuous-Time Dynamic Graph (CTDG) [99, 9]. A DTDG is similar to a spatio-temporal graph with a sequence of spatial graphs, but without spatio-temporal relationships to link the elements. A DTDG is therefore less expressive than a spatio-temporal graph. A CTDG represents evolution of a graph as a sequence of node/edge addition and deletion events. However, in existing works using graphs for SITS, objects represented by nodes are usually defined for each date independently. Thus the node set can change greatly between two acquisitions. In a CTDG framework, this can be expressed as a deletion of all previous nodes, followed by the addition of a set of new nodes. This makes it difficult to track the temporal dynamics of a phenomenon using this formulation. Therefore, DTDG and CTDG are unsuitable to model dynamic phenomena captured by SITS, thus explaining our choice to define the spatio-temporal graph as a static graph (i.e., with a fixed topology) with two types of neighborhood, one spatial and the other spatio-temporal.
@@ 3.1 Collecting the satellite image time series
In this section, the focus is given to the spatial and temporal characteristics, overlooking data storage and transmission constraints, and preprocessing such as radiometric correction [174] and geometric realignment [53]. Specifically, we can identify three common types of SITS based on their spatial resolution and revisit time, i.e., the time it takes for the satellite to acquire new data from the same location on Earth.
The first one enables a high temporal sampling (from less than an hour to one or two days), but with a low spatial resolution (hectometric at best). This kind of SITS is provided either by geostationary satellites or wide-swath satellites in low orbit to capture large, highly dynamic structures.
On the contrary, some SITS can have a high spatial resolution (from a few decimeters to several tens of meters) and a low temporal resolution, to focus on local details of long-term phenomena. Site revisits can be regular, such as once a month for low-earth orbit satellites, or irregular, if acquisitions are made on demand by an agile satellite than can point off-nadir.
Finally, acquisition settings can achieve both high spatial and temporal resolution. Adding identical satellites on the same low-Earth orbit, but delayed, forms constellations and improves revisit time without compromising spatial resolution. However, deploying mega-constellations raises the question of low-Earth orbit congestion [15].
@@
These differences in acquisition methods mean that SITS have a wide range of spatial resolutions, from a few decimeters to several kilometers, and different temporal samplings, regular or not. There is a third concern regarding data resolution. For optical SITS, it corresponds to spectral resolution ranging from panchromatic acquisition, to multispectral with a few wavelengths, often in the visible and infrared range, and hyperspectral with acquisitions over hundred of narrowed contiguous spectral bands. It should be noted that there is a trade-off between spatial and spectral resolutions, meaning that a system with high spectral resolution can only provide medium or low spatial resolution. For radar data, we can consider the number of frequency bands and the polarization of the acquisition as the radar counterpart of spectral resolution. The ability to distinguish subtle differences in measured energies is called radiometric resolution.
@@
The analysis of SITS is likely to extend beyond the multispectral and radar aspects currently studied. For example, some works manipulate multitemporal hyperspectral images in the form of time series at the pixel level. However, only a few exploit these satellite acquisitions as SITS, exploiting jointly spatial, temporal and spectral dimensions [73, 67, 125, 17, 118], certainly due to the lack of methods to process such complex and large data. Moreover, if we consider future satellite planning, the Optical Constellation in Three Dimensions (CO3D) will add depth to the spatial dimension thanks to three-dimensional acquisitions, going beyond the stereo acquisitions of Pléiades and Pléiades Neo, and thus providing multi-temporal 3D point clouds [115], which can be processed using graph-based methods.